# Beyond LLMs: An Exploration of Small Open-source Language Models in Logging Statement Generation

### v1

@@ 7. Related Work
7.1. Automated logging statement generation
Traditionally, the automation of logging statements is divided into two primary stages (Chen and Jiang, 2021; He et al., 2021): the identification of logging locations and the creation of logging statements. These stages are respectively denoted as where to log and what to log (Zhong et al., 2025). In addressing the complexities associated with determining where to log, various methodologies have been investigated by researchers to identify appropriate logging locations within source code (Jia et al., 2018; Li et al., 2020; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). Regarding what to log, the generation of logging statements is usually segmented into three specific subtasks: the generation of logging text (Ding et al., 2022), the selection of logging variables (Liu et al., 2019; Yuan et al., 2012c), and the prediction of the logging level (Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019).
@@ 7.1. Automated logging statement generation
In this paper, we distinguish our work by focusing on the use of SOLMs for automated logging statement generation, addressing the limitations of LLMs in terms of privacy, computational efficiency, and adaptability to enterprise-specific logging practices. Unlike prior proposed approaches, which predominantly rely on LLMs, our study leverages fine-tuned SOLMs. This enables local deployment, mitigating privacy risks associated with cloud-based LLM APIs, and significantly reduces computational overhead, aligning with sustainable computing goals. Furthermore, our comprehensive evaluation using the AL-Bench dataset (Tan et al., [n. d.]) demonstrates SOLMs’ robust generalization across diverse, unseen repositories, a critical aspect not extensively explored in prior work. By systematically investigating prompting strategies and fine-tuning techniques, we provide a scalable and practical solution for automated logging that balances performance with resource constraints, offering a viable alternative for real-world software development.
@@ 8. Conclusion
In this paper, we address the challenges associated with manual logging statement generation by comprehensively investigating the potential of SOLMs as a more viable alternative to resource-intensive and privacy-concerning LLMs. We conduct an extensive, large-scale empirical study to systematically evaluate the efficacy of four prominent SOLMs. This evaluation encompasses various interaction strategies, such as RAG; PEFT techniques, with a focus on LoRA; the impact of different model sizes; and the comparative performance of base versus instruction-tuned model types.
Our key findings provide several critical insights into leveraging SOLMs for this task. We demonstrate that RAG significantly enhances the performance of SOLMs in generating logging statements, and LoRA proves to be a highly effective PEFT technique, enabling substantial improvements with minimal trainable parameters. While larger SOLMs generally yield better results, this is balanced by increased computational demands, highlighting an important performance-resource trade-off. Furthermore, instruction-tuned SOLMs consistently outperform their base counterparts, benefiting from their inherent instruction-following capabilities. Most notably, our research establishes that fine-tuned SOLMs, particularly Qwen2.5-coder-14B, can surpass existing specialized logging tools and even larger LLM baselines in both the accuracy of predicting logging locations and the quality of the generated statements. These findings are further corroborated by LLM-as-a-judge evaluations, which confirm the high quality of SOLM-generated outputs. Additionally, we find that SOLMs exhibit robust generalization capabilities across diverse, unseen code repositories, underscoring their practical applicability.
In conclusion, this study provides strong evidence that appropriately fine-tuned SOLMs offer a powerful, efficient, privacy-preserving, and adaptable solution for automated logging statement generation. By making our methodology, datasets, and results publicly available (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025), we aim to stimulate further research and development in this domain, ultimately contributing to improved software maintenance practices.

### v2

@@ 7. Related Work
7.1. Automated Logging
Traditionally, the automation of logging statements is divided into two primary stages (Chen and Jiang, 2021; He et al., 2021): the identification of logging locations and the creation of logging statements. These stages are respectively denoted as where to log and what to log (Zhong et al., 2025). In addressing the complexities associated with determining where to log, various methodologies have been investigated by researchers to identify appropriate logging locations within source code (Jia et al., 2018; Li et al., 2020a; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). Regarding what to log, the generation of logging statements is usually segmented into three specific subtasks: the generation of logging text (Ding et al., 2022), the selection of logging variables (Liu et al., 2019; Yuan et al., 2012c), and the prediction of the logging level (Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019).
@@ 7.1. Automated Logging
In this paper, we distinguish our work by focusing on the use of SOLMs for automated logging, addressing the limitations of LLMs in terms of privacy, computational efficiency, and adaptability to enterprise-specific logging practices. Unlike prior proposed approaches, which predominantly rely on LLMs, our study leverages fine-tuned SOLMs. This enables local deployment, mitigating privacy risks associated with cloud-based LLM APIs, and significantly reduces computational overhead, aligning with sustainable computing goals. Furthermore, our comprehensive evaluation using the AL-Bench dataset (Tan et al., [n. d.]) demonstrates SOLMs’ robust generalization across diverse, unseen repositories, a critical aspect not extensively explored in prior work. By systematically investigating prompting strategies and fine-tuning techniques, we provide a scalable and practical solution for automated logging that balances performance with resource constraints, offering a viable alternative for real-world software development.
@@ 8. Conclusion
In this paper, we explore solutions to the challenges associated with manual logging statement generation by comprehensively investigating the potential of SOLMs as a more viable alternative to resource-intensive and privacy-concerning LLMs. We conduct an extensive, large-scale empirical study to systematically evaluate the efficacy of four prominent SOLMs. This evaluation encompasses various prompt strategies, such as RAG; PEFT techniques, with a focus on LoRA; the impact of different model sizes; and the comparative performance of base versus instruction-tuned model types.
Our key findings provide several critical insights into leveraging SOLMs for automatic logging. We demonstrate that (1) RAG significantly enhances the performance of SOLMs in generating logging statements, and (2) LoRA proves to be a highly effective PEFT technique, enabling substantial improvements with minimal trainable parameters. (3) While larger SOLMs generally yield better results, this is balanced by increased computational demands, highlighting an important performance-resource trade-off. Furthermore, (4) instruction-tuned SOLMs consistently outperform their base counterparts, benefiting from their inherent instruction-following capabilities. Most notably, (5) our research establishes that fine-tuned SOLMs, particularly Qwen2.5-coder-14B, can surpass existing specialized logging tools and even larger LLM baselines in both the accuracy of predicting logging locations and the quality of the generated statements. These findings are further corroborated by LLM-as-a-judge evaluations, which confirm the high quality of SOLM-generated outputs. Additionally, (6) we find that SOLMs exhibit robust generalization capabilities across diverse, unseen code repositories, underscoring their practical applicability.
In conclusion, this study provides strong evidence that appropriately fine-tuned SOLMs offer a powerful, efficient, privacy-preserving, and adaptable solution for automated logging. By making our methodology, datasets, and results publicly available (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025), we aim to stimulate further research and development in this domain, ultimately contributing to improved software maintenance practices.